# EVI-SAM: Robust, Real-time, Tightly-coupled Event-Visual-Inertial State Estimation and 3D Dense Mapping

### v2

@@ III-A Event Representations
The TS with polarity can encode the spatio-temporal constraints of the historical event stream at any given instant. Using an exponential decay kernel, it can emphasize recent events compared to past events. Assuming $t_{last}$ as the timestamp of the last event at each pixel coordinate $\bm{x}=(u,v)$, the TS with polarity at time $t\geq t_{last}(\bm{x})$ can be defined by:
@@
where $\Delta t$ is the constant temporal window of the observed events.
@@ IV-A Hybrid Optimization for Feature-based EVIO and Event-based Direct Tracking
where $\bm{\chi}$ is the estimated state of the system, the state is composed by the position $\bm{p}^{w}_{b}$, and the orientation quaternion $\bm{q}^{w}_{b}$, and the velocity $\bm{v}^{w}_{b}$ of the IMU in the world frame. $\bm{z}$ stands for the aggregation of the independent feature-based and direct-based measurements ($n=2$ in our framework), and $\left\{\bm{r}_{p},\bm{H}_{p}\right\}$ encapsulates the prior information or the marginalization of the system state. $\bm{r}\left(\cdot\right)$ denotes the residual function of each measurement and $||\cdot||_{\bm{P}}$ is the Mahalanobis norm. We decompose this optimization problem as two individual factors, the relative pose constraints from direct-based measurements (Eq. (8)), and the constraints from feature-based EVIO (Eq. (9)).
@@ IV-C Feature-based EVIO Pose Tracking
Eq. (9) is composed of three components: (i) the IMU pre-integration residuals $\bm{e}^{k}_{imu}$; (ii) the feature-based image residuals $\bm{e}^{k,l}_{image}$; (iii) the feature-based event residuals $\bm{e}^{k,l}_{event}$. $\zeta$ and $\xi$ are the set of event-corner features and image features, respectively. $K$ ($K=10$ in our experiments) is the total number of keyframes in the sliding window. The detailed mathematical explanation can be found in our previous work [15]. Here, we only review the re-projection constraint of the event-corner features. Considering the $l^{th}$ event-corner feature that is first observed in the $i^{th}$ keyframe, the residual for its observation in the $k^{th}$ keyframe is defined as:
@@ VI Evaluation
Secondly, we assess the mapping performance of EVI-SAM through four sets of experiments. The first set (Section VI-B) evaluates the mapping performance across diverse scenarios to show the generation ability of our event-based dense mapping method. Section VI-C compares our 3D mapping approach with other 3D reconstruction baselines. Moreover, Section VI-D evaluates our mapping performance under challenging situations, such as HDR (Section VI-D1), and the scenarios involving aggressive motion (Section VI-D2). Section VI-E evaluates the overall system on the onboard platform, including comparing local mapping with a commercial depth camera and assessing the global event-based 3D dense reconstruction performance.
@@ VI-E1 Event-based Handheld Device
Our handheld device is shown in Fig. 11, which includes a power supply unit, an onboard computer NUC (equipped with Intel i7-1260P, 32GB RAM, and Ubuntu 20.04 operation system), a DAVIS346 event camera, and a Intel® RealSense™ D455 RGB-D camera. All mechanical modules of this device are designed for 3D printing, and both design schematics and the collected data sequence are available on our GitHub repository: https://github.com/arclab-hku/Event_based_VO-VIO-SLAM?tab=readme-ov-file#5-evi-sam
VI-E2 Local Mapping Performance
We conduct real-time testing of EVI-SAM using the handheld device within the LG01_lab of the Haking Wong Building at The University of Hong Kong. The results, including event-based dense mapping at selected viewpoints and the estimated trajectory, are illustrated in Fig. 13. Three live demonstrations are also available at https://b23.tv/cwetpxL. Additionally, we conduct qualitative comparisons between the dense mapping results from our EVI-SAM and the depth images from the RGB-D camera. To ensure a fair comparison, we render the texture information from the intensity image of DAVIS346 onto the depth image of the RGB-D camera, maintaining the same resolution as the DAVIS346. The depth images are directly acquired from the RGB-D camera, with its infrared ranging sensor obscured. We directly align the texture point clouds from the RGB-D camera along with the estimated pose of our tracking module. As observed in the highlighted area within the red circle in Fig. 12, it is evident that our proposed event-based dense mapping approach demonstrates performance comparable to commercial depth cameras. The reconstructed depths from the RGB-D camera are notably noisy and cluttered, lagging far behind the results obtained from our EVI-SAM. Moreover, it exhibits superior noise reduction and texture recovery capabilities.
@@ VI-G Limitations and Discussions
Certainly, image-based dense reconstruction has been thoroughly studied and achieved promising results. However, the exceptional advantages of event cameras underscore the importance of exploring event-based dense reconstruction, especially considering that non-learning methods for event-based dense reconstruction remain an unexplored territory. Our method generates dense depths from event streams, with images only serving as guidance. Thus the reconstruction results may not match the impressive performance of some state-of-the-art RGB-only or RGB-D methods. Nevertheless, our focus lies in addressing dense reconstruction using event cameras, which offer a potential solution for challenging scenarios not adequately addressed by standard images. This work not only aims to achieve optimal mapping performance in challenging situations but also serves as an inspiration for the research community encouraging future research in event-based dense reconstruction.

### v3

@@ III-A Event Representations
The TS with polarity can encode the spatio-temporal constraints of the historical event stream at any given instant. Using an exponential decay kernel, it can emphasize recent events compared to past events. Assuming $t_{last}(\bm{x})$ as the timestamp of the last event at each pixel coordinate $\bm{x}=(u,v)$, the TS with polarity at time $t\geq t_{last}(\bm{x})$ can be defined by:
@@
where $\Delta t$ is the constant temporal window of the observed events. While the temporal window length for event processing remains consistent with the frequency of the input event stream.
III-B Initialization
The event-based mapping thread and the direct event-based tracking are initiated after the successful bootstrap of the feature-based EVIO. Adopting from [45, 46], the feature-based EVIO pipeline commences with a vision-only structure from motion (SfM) to establish the up-to-scale structure of camera pose and event-corner feature positions. By loosely aligning the SfM with the pre-integrated IMU measurements, it can bootstrap the system from unknown initial states.
@@ IV-A Hybrid Optimization for Feature-based EVIO and Event-based Direct Tracking
where $\bm{\chi}$ is the estimated state of the system, the state is composed by the position $\bm{p}^{w}_{b}$, and the orientation quaternion $\bm{q}^{w}_{b}$, and the velocity $\bm{v}^{w}_{b}$ of the IMU in the world frame. $\bm{z}$ stands for the aggregation of the independent feature-based and direct-based measurements ($n=2$ in our framework), and $\left\{\bm{r}_{p},\bm{H}_{p}\right\}$ encapsulates the prior information or the marginalization [51, 52] of the system state. $\bm{r}\left(\cdot\right)$ denotes the residual function of each measurement and $||\cdot||_{\bm{P}}$ is the Mahalanobis norm. We decompose this optimization problem as two individual factors, the relative pose constraints from direct-based measurements (Eq. (8)), and the constraints from feature-based EVIO (Eq. (9)).
@@ IV-C Feature-based EVIO Pose Tracking
Eq. (9) is composed of three components: (i) the IMU pre-integration residuals $\bm{e}^{k}_{imu}$; (ii) the feature-based image residuals $\bm{e}^{k,l}_{image}$; (iii) the feature-based event residuals $\bm{e}^{k,l}_{event}$. $\zeta$ and $\xi$ are the set of event-corner features and image features, respectively. $K$ ($K=10$ in our experiments) is the total number of keyframes in the sliding window. The detailed mathematical explanation and the keyframe selection scheme can be found in our previous work [15]. Here, we only review the re-projection constraint of the event-corner features. Considering the $l^{th}$ event-corner feature that is first observed in the $i^{th}$ keyframe, the residual for its observation in the $k^{th}$ keyframe is defined as:
@@ VI Evaluation
Secondly, we assess the mapping performance of EVI-SAM through four sets of experiments. The first set (Section VI-B) evaluates the mapping performance across diverse scenarios to show the generation ability of our event-based dense mapping method. The second set of experiments compare our 3D mapping approach with different mapping baselines, including those relying on event-based (Section VI-C), image-based (APPENDIX A-A), learning-based (APPENDIX A-B), and NeRF-based (APPENDIX A-C) methods. Moreover, Section VI-D evaluates our mapping performance under challenging situations, such as HDR (Section VI-D1), and the scenarios involving aggressive motion (Section VI-D2). Section VI-E evaluates the overall system on the onboard platform, including comparing local mapping with a commercial depth camera and assessing the global event-based 3D dense reconstruction performance.
@@ VI-C2 Quantitative Comparison
Event-only stereo failed failed 0.26 / 100%
Ref. [10] Image-only stereo failed failed 0.24 / 100%
Event-Image stereo failed failed 0.22 / 100%
@@
Furthermore, Table III also provides the depth error of a learning-based method [10]. Since we followed the same evaluation protocol as Ref. [10], we can directly report the raw results from the original paper. Despite this method being numerically superior to the other non-learning methods, it requires scene-specific pre-training for each data sequence and high-level hardware resources such as GPU. It is worth noting that when evaluating the indoor_flying3, both indoor_flying1 and indoor_flying2 sequences are used to train the network. Additionally, only around 1000 frames from the entire rosbag, containing over 3000 frames, can be used for testing. In contrast, our EVI-SAM can process the entire rosbag of indoor_flying1 to indoor_flying3 without any pre-training. Therefore, we only report the raw result of indoor_flying3 from Ref. [10], and label the result of indoor_flying1 and indoor_flying2 as failed since it only works under specified pre-training conditions. The qualitative comparison with this learning-based method is available in APPENDIX A-B.
@@ VI-E1 Event-based Handheld Device
Our handheld device is shown in Fig. 11, which includes a power supply unit, an onboard computer NUC (equipped with Intel i7-1260P, 32GB RAM, and Ubuntu 20.04 operation system), a DAVIS346 event camera, and a Intel® RealSense™ D455 RGB-D camera. All mechanical modules of this device are designed for 3D printing, and both design schematics and the collected data sequence are available on our GitHub repository:https://github.com/arclab-hku/Event_based_VO-VIO-SLAM/blob/main/EVI-SAM/data_srcipt.md
@@
VI-E2 Local Mapping Performance
We conduct real-time testing of EVI-SAM using the handheld device within the LG01_lab of the Haking Wong Building at The University of Hong Kong. The results, including event-based dense mapping at selected viewpoints and the estimated trajectory, are illustrated in Fig. 13. Three live demonstrations are also available at https://b23.tv/cwetpxL. Additionally, we conduct qualitative comparisons between the dense mapping results from our EVI-SAM and the depth images from the RGB-D camera. To ensure a fair comparison, we render the texture information from the intensity image of DAVIS346 onto the depth image of the RGB-D camera, maintaining the same resolution as the DAVIS346. The depth images are directly acquired from the RGB-D camera, with its infrared ranging sensor obscured. We directly align the texture point clouds from the RGB-D camera along with the estimated pose of our tracking module. As observed in the highlighted area within the red circle in Fig. 12, it is evident that our proposed event-based dense mapping approach demonstrates performance comparable to commercial depth cameras. The reconstructed depths from the RGB-D camera are notably noisy and cluttered, lagging far behind the results obtained from our EVI-SAM. Moreover, it exhibits superior noise reduction and texture recovery capabilities. More dense mapping performance comparisons between our EVI-SAM and the RGB-D camera can be found in the APPENDIX A-A.
@@ VI-G Limitations and Discussions
Certainly, image-based dense reconstruction has been thoroughly studied and achieved promising results. However, the exceptional advantages of event cameras underscore the importance of exploring event-based dense reconstruction, especially considering that non-learning methods for event-based dense reconstruction remain an unexplored territory. Our method generates dense depths from event streams, with images only serving as guidance. Thus the reconstruction results may not match the impressive performance of some state-of-the-art RGB-only or RGB-D methods. Besides, our method does not rely on images to generate dense depth, instead, dense depths are generated from the event stream with images only serving as guidance. Therefore, the reconstruction results may not be as impressive as some state-of-the-art RGB-based methods. Nevertheless, our focus lies in addressing dense reconstruction using event cameras, which offer a potential solution for challenging scenarios not adequately addressed by standard images. This work not only aims to achieve optimal mapping performance in challenging situations but also serves as an inspiration for the research community encouraging future research in event-based dense reconstruction.
@@ VII CONCLUSIONS
Appendix A More Mapping Performance Comparisons
Section VI-C provides a quantitative and qualitative comparison of our event-based dense mapping performance with other event-based mapping baselines. This APPENDIX further extends the mapping performance comparison with traditional image-based methods (APPENDIX A-A), learning-based mapping method (APPENDIX A-B), and NeRF-based methods (APPENDIX A-C), respectively.
Figure A-1: Local mapping performance comparison. The first column shows the intensity image of the selected view. Column 2 to 5 show the color point cloud results from our EVI-SAM, stereo RGB camera [73], RGB-D camera [74], and monocular camera [57], respectively.
Figure A-2: Qualitative comparison of our EVI-SAM and the learning-based method [10] in MVSEC [66] dataset. For each element, the first row is the (i) image view, and the estimated depth of learning-based method, which utilizes the (ii) event-only, the (iii) image-only, and the (iv) event-image methods; The second row is the (v) ground truth depth obtained through 3D LiDAR, the estimated depth of our EVI-SAM using (vi) event-only and (vii) event-image, and our estimated depth with texture information.
A-A Mapping Performance Comparison with Traditional Image-based Methods
We conduct qualitative comparisons of the depth estimation results presented as texture point clouds obtained from our EVI-SAM, with the depth estimation results from stereo RGB camera, monocular RGB camera, and RGB-D camera serving as the baseline. We used the data sequence ”LG_office”, which features fast motion, from Section VI-E1. To ensure a fair comparison, firstly, we only employ lightweight and real-time methods to estimate depth from standard cameras, as opposed to offline or complex approaches that rely on GPU acceleration. Since our EVI-SAM operates in real-time solely on the CPU. Secondly, we resize the resolution of the depth generated from the baselines to match the resolution of our event camera (346 $\times$ 260). Thirdly, we render the texture information from the intensity image of the DAVIS346 onto the estimated depth to generate the textured point cloud. More details these baseline methods are as follows:
  • RGB-D camera. The results are directly obtained from the depth image of Realsense D455 [74], with its infrared ranging sensor obscured. While the texture information is transformed from the intensity image of the DAVIS346.
  • Stereo RGB cameras. The results are generated by the image_undistort tool [73].
  • Monocular RGB camera. The results are generated by using the monocular dense mapping algorithm in Ref. [57] using GPU acceleration for real-time running.
  • All of these baseline methods use the monocular image+IMU (VINS-MONO [45]) to provide pose estimation.
As can be seen from Fig. A-1, our EVI-SAM performs better than these baselines in both the accuracy of depth recovery and the integrity of texture information. The textured point cloud obtained from the RGB-D camera is directly captured from a commercial Realsense camera. However, its infrared ranging sensor is obstructed, as infrared light can interfere with the perception of event cameras. When the infrared ranging sensor is blocked, the depth map produced by the Realsense camera tends to contain a considerable amount of noise, as illustrated in column 4 of Fig. A-1. This observation matches our expectations, as demonstrated in Section VI-E2, where it is evident that the depth map generated by the RGB-D camera exhibits more noise compared to that of our EVI-SAM.
Regarding the results from monocular and stereo RGB cameras, as shown in columns 3 and 5 of Fig. A-1, although their estimated depths exhibit less noise compared to those estimated by RGB-D cameras, the completeness of depth recovery falls far behind that of RGB-D and our EVI-SAM. This might be caused by the challenging testing conditions characterized by varying illumination and fast motion. Additionally, the regions with limited texture and the drawbacks such as the narrow dynamic range of image sensor perception lead to the poor depth recovery performance of monocular and stereo RGB cameras on this data sequence. While our EVI-SAM still can provide reliable depth perception under such challenging situations.
A-B Mapping Performance Comparison with Learning-based Mapping Method
In Section VI-C2, we conduct the quantitative comparisons with learning-based approach [10]. As shown in Fig. A-2, we further demonstrate the qualitative comparisons with this learning-based approach using the sequences of indoor_flying3 from the MVSEC dataset [66]. Since the source code and pre-training model of Ref. [10] are not released, the estimated depth results are borrowed from the original paper. While the ground truth depth and our estimated depth are pseudo-colored, from red (close) to blue (far). For comparison, we select the reference views that are similar to those in Ref. [10]. It is worth noting that this learning-based approach employs a complex network with many structures, such as recycling, deformable, and multiscale architecture. Although its numerical results in Table III may appear superior to our EVI-SAM, a thorough examination of the qualitative results indicates otherwise. There are instances of significant noise in their estimated results, particularly when employing stereo images or stereo event setups. Conversely, our EVI-SAM not only demonstrates superior mapping accuracy (both quantitatively and qualitatively) but also can operate without any scene-specific pre-training.
Figure A-3: Qualitative comparison of our EVI-SAM, the Instant-NGP [75], and the ground truth image view. The red box highlights the difference of the reconstruction quality and the motion blur caused by image blur in NeRF-based methods. While our event-based dense mapping demonstrates performance comparable to state-of-the-art NeRF-based mapping.
Additionally, our EVI-SAM consistently achieves good mapping results across diverse and challenging scenarios, as evidenced in Sections VI-B and VI-D. In particular, thorough the evaluations in Section VI-D2, our event-based dense mapping reliably reconstructs the 3D structure of scenes even under fast rotation (up to 290°$\degree/s$) and translation (up to 18 $m^{2}/s$) motions. While most of the event-based mapping works, especially for those relying on learning-based methods, often neglect to assess their performance under aggressive motion.
A-C Mapping Performance Comparison with NeRF-based Methods
We evaluate the NeRF-based methods using the LG_office sequence of the EVI-SAM dataset in Section VI-E1. Neural radiance field (NeRF) is a fully-connected neural network that can generate novel views of complex 3D scenes, based on a partial set of 2D images. It operates by processing input images representing a scene and interpolating between them to render one complete scene. Therefore, each scene must be retrained and can only be tested in the same scene. Failure to retrain the model when the scene changes would result in entirely incorrect outputs. We employ the Instant-NGP [75] for the novel view synthesis from the image of DAVIS346 and use the COLMAP [76] to estimate the camera parameters, including both intrinsics and extrinsics (pose), for each input image. For comparing the performance of our event-based dense mapping, we visualize the view synthesis of similar viewpoints of Fig. 13 and Fig. A-1.
As highlighted with red-box in Fig. A-3, significant blurring effects can be observed in the reconstruction results of the NeRF-based method, which is caused by image blurring in rapid motion. In contrast, our event-based dense mapping method effortlessly overcomes this challenge without encountering motion blur. Additionally, the memory consumption of Instant-NGP is extremely high, making it impossible for our GPU (NVIDIA GeForce RTX 4070 Ti) to complete training with the entire sequence. Therefore, we had to select 8 corresponding viewpoints for 8 separate training sessions. In other words, the result of the Instan-NGP in Fig. A-3 stems from 8 entirely distinct training sessions, whereas our EVI-SAM can operate in real-time without the need for segmentation processing or prior training.
Furthermore, we also utilized the RGB-D data from sequence to train two state-of-the-art NeRF-based RGB-D SLAM [77, 78]. However, due to significant noise in the depth camera and the lack of an infrared ranging sensor, depth information cannot be captured in low-texture scenes (i.e. the white wall in Fig. A-3). This resulted in inadequate depth information for training both Nice-SLAM [77] and Co-SLAM [78]. Besides, the rapid and aggressive motion in the testing sequence scenes is too challenging for these methods, resulting in inadequate pose estimation and ultimately leading to mapping failures. The mapping results of Co-SLAM are shown in Fig. A-4.
Figure A-4: The mapping result of Co-SLAM [78] and the corresponding selected viewpoint.
Appendix B Ablation Study of the Pose Tracking
B-A Ablation Study of Event and Image-based VIO
As indicated in Section VI-A, compared to standard cameras, event cameras are capable of providing reliable visual perception during high-speed motion and HDR scenarios. However, when event cameras have restricted relative motion, such as in static states, they may produce limited information. Although the image camera encounters difficulties under high-speed motion or HDR scenarios, it can provide rich-intensity textures of the scenes under uniform motion or favorable lighting conditions. Therefore, leveraging the complementary advantages of both standard and event cameras is significant and promising, as emphasized in many previous works [15, 16, 14, 10, 38, 13, 62, 20].
Although the effectiveness of tightly integrating events, images, and IMU for pose estimation has been thoroughly assessed in various event-based studies, we still conduct this ablation study of various combinations within our EVI-SAM framework including event-IMU-based (EIO), image-IMU-based (VIO), and event-image-IMU-based (EVIO) configurations, as shown in Table B-I. In the first sequence, our feature-based EIO and VIO demonstrate comparable performance, thanks to our well-designed framework and the HDR level is not serious. Conversely, in the second sequence, our feature-based EIO significantly outperforms the feature-based VIO, attributed to the high motion intensity in the testing scene. While our featured-based EVIO performs satisfactory results and surpasses the solely image-based VIO in both testing sequences.
TABLE B-I: Accuracy Comparison of various combinations within EVI-SAM framework, including Event+IMU, Image+IMU, Event+Image+IMU, direct-based, and our EVI-SAM
Sequence Feature-based Direct-based Hybrid
EIO VIO EVIO EVIO EVIO
vicon_hdr4 0.115 0.116 0.089 0.080 0.092
vicon_aggressive_hdr 0.572 1.266 0.921 0.594 0.683
  • Unit:%, 0.115 means the average error would be 0.115m for 100m motion; Aligning the whole ground truth trajectory with estimated poses.
B-B Ablation Study of Direct-based and Feature-based EVIO
Section VI-A provides the analysis of the performance comparison between feature-based and direct-based EIO/EVIO, both in the same framework (Fig. 6(a)6(b)) and across different frameworks (Fig. 6(c)6(d), and Table III). In this APPNDIX, we further demonstrates more performance comparisons among different combinations within the same framework (our EVI-SAM). We used the publicly available monocular HKU-Dataset [22] (346$\times$260, event, image, IMU, GT pose) and the DAVIS240C Dataset [59] (240$\times$180, event, image, IMU, GT pose). Since the DAVIS240C dataset typically exhibits limited displacement in the last 30 seconds, this is not conducive to explicitly evaluating the differences in position accuracy. Therefore, we edited the rosbag files to include only the [0-30s] segment of the total duration of the 60s. The sequences in Table B-II are renamed with ”edited” to distinguish them from the complete data in Table I.
TABLE B-II: Accuracy comparison of various combinations within EVI-SAM framework, including feature-based, direct-based, and our hybrid (feature-based + direct-based) tracking pipeline
Sequence Feature-based Direct-based Hybrid EVI-SAM
vicon_hdr4 0.09 0.08 0.09
vicon_aggressive_hdr 0.92 0.59 0.68
boxes_translation_edited 0.15 0.10 0.14
hdr_boxes_edited 0.21 0.12 0.21
boxes_6dof_edited 0.34 0.16 0.20
dynamic_translation_edited 0.40 0.32 0.34
dynamic_6dof_edited 0.19 0.18 0.23
poster_translation_edited 0.12 0.15 0.35
hdr_poster_edited 0.55 0.32 0.44
poster_6dof_edited 0.10 0.08 0.09
  • Unit:%, 0.10 means the average error would be 0.10m for 100m motion; Aligning the whole ground truth trajectory with estimated poses.
As can be seen from the result, the proposed direct-based pose tacking outperforms the feature-based component in most of the sequences. This result is consistent with the theoretical analysis in Section IV-A, as well as the common viewpoint in direct-based methods, i.e. sub-pixel alignment of the direct-based methods normally achieves better accuracy than the feature-based methods. On the other hand, as what has been well evaluated in our previous works [15, 16], our feature-based EVIO can operate robustly in large-scale outdoor environments and flight scenarios. This proves that it is suitable for handling irregular scene variations and significant inter-frame motions. Therefore, leveraging the complementary advantages of direct-based and feature-based methods is significant and promising.
B-C More Tracking Performance Comparison in Challenge Situations
Section VI-A provides the performance comparison of our event-based hybrid pose tracking with the state-of-the-art EIO, VIO, and EVIO works in challenging situations. We further conduct the comparison of our EVI-SAM with two latest works in the field of image-based VIO [50, 79]. We used the publicly available stereo HKU-dataset [16] which features rapid motion and HDR scenarios. Following the trajectory alignment protocol of Table II, we computed the mean position as percentages of the total traveled distance, while the estimated trajectories and ground-truth were aligned in SE3 with all alignments.
TABLE B-III: Accuracy Comparison of Our EVI-SAM with image-based VIO on HKU-Dataset
Sequence EnVIO [50] Stereo VIO MSOC-S-IKF [79] Stereo VIO MSOC-S-IKF [79] Mono VIO Our EVI-SAM Mono EVIO
hku_agg_translation failed failed 0.51 0.17
hku_agg_rotation 2.12 1.52 3.50 0.24
hku_agg_flip 2.94 failed failed 0.32
hku_agg_walk 3.38 failed 2.07 0.26
hku_hdr_circle 0.85 failed failed 0.13
hku_hdr_slow 0.43 failed failed 0.11
hku_hdr_tran_rota 0.37 failed 5.48 0.11
hku_hdr_agg 0.50 failed failed 0.10
hku_dark_normal failed failed failed 0.85
  • Unit:%, 0.17 means the average error would be 0.17m for 100m motion; Aligning the whole ground truth trajectory with estimated poses.
Table B-III demonstrates that our EVI-SAM outperforms these image-based VIO methods in all sequences. Given the challenge of these sequences, which involve rapid movement, low-light conditions, and significant illumination change, it is difficult for the traditional image-based methods [63, 64, 50, 79] to handle such high-speed motions and HDR scenarios. For instance, both monocular and stereo versions of the MSOC-S-IKF [79] perform well at the beginning of most of these sequences. However, when confronted with fast motion or sudden illumination changes, these VIO systems are prone to failure or significant drift due to the unreliable perception of images. Certainly, the performance may vary among different image-based VIO methods. For example, although the results of ORB-SLAM3 [63] in Table II and the EnVIO [50] also exhibit large errors under such challenging conditions, they tend to avoid serious drift or failure, possibly owing to differences in framework stability.